# Detecting LLM Hallucination Through Layer-wise Information Deficiency: Analysis of Unanswerable Questions and Ambiguous Prompts

### v1

@@ Abstract
Large language models (LLMs) frequently generate confident yet inaccurate responses, introducing significant risks for deployment in safety-critical domains. We present a novel approach to detecting model hallucination through systematic analysis of information flow across model layers when processing inputs with insufficient or ambiguous context. Our investigation reveals that hallucination manifests as usable information deficiencies in inter-layer transmissions. While existing approaches primarily focus on final-layer output analysis, we demonstrate that tracking cross-layer information dynamics ($\mathcal{L}$I) provides robust indicators of model reliability, accounting for both information gain and loss during computation. $\mathcal{L}$I improves model reliability by immediately integrating with universal LLMs without additional training or architectural modifications.
Detecting LLM Hallucination Through Layer-wise Information Deficiency: Analysis of Unanswerable Questions and Ambiguous Prompts
Hazel Kim, Adel Bibi, Philip Torr, Yarin Gal University of Oxford {hazel.kim, yarin.gal}@cs.ox.ac.uk, {adel.bibi, philip.torr}@eng.ox.ac.uk
$\dagger$\$$\dagger$\$footnotetext: Correspondence to Hazel Kim <hazel.kim@cs.ox.ac.uk>.$\ddagger$\$$\ddagger$\$footnotetext: Equal advising.
@@ 1 Introduction
Large language models (LLMs) have achieved unprecedented success across diverse natural language tasks, particularly in complex reasoning ranging from commonsense to arithmetic knowledge (Achiam et al., 2023; Touvron et al., 2023; Abdin et al., 2024). However, these models face a critical challenge known as hallucination—a phenomenon where responses appear convincingly authoritative despite being inaccurate (Ji et al., 2023; Xu et al., 2024b; Liu et al., 2023). While numerous empirical studies have investigated potential sources of hallucination, recent theoretical work by Xu et al. (2024a) demonstrates the fundamental impossibility of eliminating this issue through any computable function.
@@
Figure 1: Our proposed $\mathcal{L}$I correlates more strongly with question difficulty as influenced by answerability and prompt clarity—compared to final layer entropy (i.e., $\mathcal{V}$-usable information). The AUROC shows model classification in (un)answerable questions. ’With prompts’ are binary-type instruction prompts appended to the question (see Table 1).
(a) $\mathcal{V}$-usable information ($\mathcal{V}$I) per each layer. The information is not monotonically increased or decreased as the layer goes deeper. The original estimation of $\mathcal{V}$-usable information applies to the final layer (i.e., final layer $\mathcal{V}$I in Fig. 1) as illustrated in red.
(b) Cumulative $\mathcal{V}$I from the beginning to the layer $i$. The final layer tends to increase the overall scores, and sometimes exceeds those of others. It is crucial to cumulate the quantities up to the final layer to represent $\mathcal{L}$I, shown as the last quantities on x-axis. Relevant cases are further illustrated in Fig. 6.
Figure 2: $\mathcal{L}$I Estimation. The reason we consider the overall layer information rather than just the final layer is that it provides a more comprehensive understanding of the model’s behavior. The final layer $\mathcal{V}$I shows almost contradictory results from $\mathcal{L}$I; the CoQA has the lowest final layer $\mathcal{V}$I and the highest $\mathcal{L}$I, for example.
@@
Prior work on $\mathcal{V}$-usable information established that computation can create usable information during feature extraction, constituting a violation of the data processing inequality (DPI) in information theory (LeCun et al., 2015; Xu et al., 2020). Our research extends this understanding by demonstrating that LLMs both create and lose usable information during layer-wise updates. As illustrated in Fig.2(a), information flow is non-monotonic across layers, highlighting the limitation of analyzing only the final layer’s computation. The proposed $\mathcal{L}$I framework captures this dual nature of DPI compliance and violation inherent to LLMs, with Fig.2 revealing substantial divergence between $\mathcal{L}$I and $\mathcal{V}$-usable information patterns. An additional explanation of DPI is presented in Section 3.2.
Prior work by Hewitt et al. (2021) has extended probing techniques on layers. However, their investigation remains limited in scope, primarily comparing layer-wise information against baseline performance and exploring part-of-speech information within the models.
To evaluate our framework, we focus on scenarios where LLMs must respond to queries with constrained contextual information (i.e., unanswerable questions), a setting particularly prone to hallucination. Our case study in Table 1 demonstrates that $\mathcal{L}$I strongly correlates with question difficulty—as influenced by answerability and prompt clarity—compared to $\mathcal{V}$I, which shows no significant correlation with model predictions.

### v2

# Detecting LLM Hallucination Through Layer-wise Information Deficiency: Analysis of Ambiguous Prompts and Unanswerable Questions

@@ Abstract
Large language models (LLMs) frequently generate confident yet inaccurate responses, introducing significant risks for deployment in safety-critical domains. We present a novel, test-time approach to detecting model hallucination through systematic analysis of information flow across model layers. We target cases when LLMs process inputs with ambiguous or insufficient context. Our investigation reveals that hallucination manifests as usable information deficiencies in inter-layer transmissions. While existing approaches primarily focus on final-layer output analysis, we demonstrate that tracking cross-layer information dynamics ($\mathcal{L}$I) provides robust indicators of model reliability, accounting for both information gain and loss during computation. $\mathcal{L}$I integrates easily with pretrained LLMs without requiring additional training or architectural modifications.
Detecting LLM Hallucination Through Layer-wise Information Deficiency: Analysis of Ambiguous Prompts and Unanswerable Questions
Hazel Kim, Tom A. Lamb, Adel Bibi, Philip Torr, Yarin Gal University of Oxford {hazel.kim, yarin.gal}@cs.ox.ac.uk, {thomas.lamb, adel.bibi, philip.torr}@eng.ox.ac.uk
[…]
@@ 1 Introduction
Figure 1: Distribution of $\mathcal{V}$-information ($\mathcal{V}$I) values in first and last layers, and $\mathcal{L}$-information ($\mathcal{L}$I) values (summation of $\mathcal{V}$I scores across all layers), as a function of prompt ambiguity. Results compare two prompt categories: (1) no instruction prompts and (2) binary instruction prompts (’Is this answerable?’).
Large language models (LLMs) have achieved unprecedented success across diverse natural language tasks, particularly in complex reasoning ranging from commonsense to arithmetic knowledge (Achiam et al., 2023; Touvron et al., 2023; Abdin et al., 2024). However, these models face a critical challenge known as hallucination, a phenomenon where responses appear convincingly authoritative despite being inaccurate (Ji et al., 2023; Xu et al., 2024b; Liu et al., 2023). While numerous empirical studies have investigated potential sources of hallucination, recent theoretical work by Xu et al. (2024a) demonstrates the fundamental impossibility of eliminating this issue through any computable function.
@@
Figure 2: $\mathcal{V}$-usable information ($\mathcal{V}$I; Xu et al. (2020)) across layers for CondaQA, CoQA, and QuAC. Solid lines denote models without instruction prompts and dashed lines denote models with binary prompts ("Is this answerable?"). The quantity of $\mathcal{V}$I is not monotonically increased or decreased across depth. By the final layer (32), the relative effect of prompts becomes inconsistent across datasets, highlighting the value of analyzing all layers.
@@
Prior work on $\mathcal{V}$-usable information established that computation can create usable information during feature extraction, constituting a violation of the data processing inequality (DPI) in information theory (LeCun et al., 2015; Xu et al., 2020). Our research extends this understanding by demonstrating that LLMs both create and lose usable information during layer-wise updates. As illustrated in Fig.2, information flow is non-monotonic across layers, highlighting the limitation of analyzing only the final layer’s computation.
To evaluate our framework, we focus on scenarios where LLMs must respond to queries with instruction prompts of different ambiguities (i.e., ambiguous prompts) and with constrained contextual information (i.e., unanswerable questions), settings particularly prone to hallucination. Our case study in Table 1 demonstrates that $\mathcal{L}$I strongly correlates with the difficulty of the question, influenced by the answerability and the clarity of the prompt, unlike $\mathcal{V}$I that shows no significant correlation with the model prediction.
[…]
Table 1: The $\mathcal{L}$I scores provide a more comprehensive overview of prompt ambiguity compared to $\mathcal{V}$I. Prediction: prediction generated by a language model — ✓: correct, ✗: incorrect. $\mathcal{V}$I: $\mathcal{V}$-usable information only applied to the final layer (Xu et al., 2020). $\mathcal{L}$I: Layer-wise usable information accumulated across layers (our proposed method).